# Complex Reasoning over Logical Queries on Commonsense Knowledge Graphs

### v1

@@ Abstract
Event commonsense reasoning requires the ability to reason about the relationship between events, as well as infer implicit context underlying that relationship. However, data scarcity makes it challenging for language models to learn to generate commonsense inferences for contexts and questions involving interactions between complex events. To address this demand, we present Com (COMplex COMmonsense), a new dataset created by sampling multi-hop logical queries (e.g., the joint effect or cause of both event A and B, or the effect of the effect of event C) from an existing commonsense knowledge graph (CSKG), and verbalizing them using handcrafted rules and large language models into multiple-choice and text generation questions.
Our experiments show that language models trained on Com exhibit significant improvements in complex reasoning ability, resulting in enhanced zero-shot performance in both in-domain and out-of-domain tasks for question answering and generative commonsense reasoning, without expensive human annotations.111Code and data are available at https://github.com/tqfang/complex-commonsense-reasoning
@@ 1 Introduction
In this paper, we construct Com (COMplex COMmonsense), a novel commonsense reasoning dataset using multi-hop queries in commonsense knowledge graphs to construct question answer pairs requiring complex narrative reasoning. To build this dataset, we use conjunctive logical queries Hamilton et al. (2018), a subset of First-Order Logical queries that use existential quantifiers and conjunction. The multi-hop projection operation involves inferring hidden contexts, while the intersection operation enables reasoning among multiple events, encompassing common cause or effect, and abduction. For example, in Figure 1, an intersection of two triples can be verbalized to a short narrative, and the process of inferring the common tail can be seen as an abduction of the hidden cause between the two heads.
To address the challenges above, we propose to first densify the CSKG to merge nodes with high semantic similarity, increasing the connectivity of the graph. Then, we use an off-the-shelf plausibility scorer to filter out low quality triples, avoiding error propagation as we construct more complicated queries. Finally, we verbalize the queries to a natural language context with handcrafted rules and Large Language Models to derive coherent and informative narrative contexts for our questions. Our final Com dataset comprises 790K question-answer pairs (both with multiple-choice and generative answer settings), including 1.3K examples that we manually verify for evaluation.
Our results demonstrate the challenges faced by even powerful LLMs and supervised question answering models on the Com dataset, underscoring the difficulty of performing complex multi-hop reasoning. Moreover, fine-tuning question answering models and generative commonsense inference models on Com leads to substantial improvements across eight commonsense reasoning datasets, showing the efficacy of our framework for boosting commonsense reasoning ability.
To conclude, our contributions are three-fold. First, we present a pipeline for sampling and verbalizing complex logical queries from CSKGs, to form a complex commonsense reasoning benchmark, Com, with minimal human effort. Second, we benchmark the complex reasoning ability of various state-of-the-art language models and question answering models on Com. Finally, we validate the benefit of fine-tuning on Com on eight zero-shot commonsense reasoning datasets.
@@ 2 Background and Related Work
Recent advances in commonsense reasoning have been driven by the construction of human-annotated (Speer et al., 2017; Sap et al., 2019a; Hwang et al., 2021; Jiang et al., 2021; Mostafazadeh et al., 2020) and human-validated (West et al., 2022; Gao et al., 2023) CommonSense Knowledge Graphs (CSKG). A common approach to create challenges for commonsense reasoning involves constructing tasks in the form of question-answering Talmor et al. (2019); Sap et al. (2019b), knowledge base completion Malaviya et al. (2020); Yang et al. (2023) and population Fang et al. (2021b, a), grounding Gao et al. (2022), and daily dialogue Kim et al. (2023), based on CSKGs. However, most of those previous benchmarks are based on one-hop triples.
@@ 3 Methodology
In this section, we introduce the construction details of Com, including the pre-processing, sampling, and verbalization of complex queries, as well as the details of human annotations.
@@ 3.1 Pre-processing
CSKGs are usually highly sparse compared to factual KGs due to the diversity and scale of commonsense Malaviya et al. (2020), resulting in many isolated nodes that can hardly be sampled as part of a complex query. To alleviate this issue, we develop a set of rules and use sentence embedding similarity to merge nodes in the CSKG, leading to 22.4% of nodes being merged and an average degree increase of 25.3%.
@@ 3.3 Verbalization
| Fine-tuned on Com | |||||||
| DeBERTa-v3-Large (+Com) | 60.09 | 58.36 | 69.41 | 61.86 | 59.29 | 81.48 | 62.79 |
| CAR-DeBERTa-v3-Large (+Com) | 61.22 | 56.13 | 69.86 | 68.56 | 56.64 | 85.19 | 63.78 |
@@
| a-NLI | CSQA | PIQA | SIQA | WG | Avg. | Com | ||
@@
| CAR Wang et al. (2023a) | ATM | 79.6 | 69.3 | 78.6 | 64.0 | 78.2 | 73.9 | 39.8 |
| HyKAS + Com(Ours) | ATM, Com | 78.4 | 69.9 | 78.7 | 64.1 | 78.3 | 73.9 | 62.8 |
| CAR + Com(Ours) | ATM Com | 81.2 | 70.9 | 80.3 | 65.6 | 77.4 | 75.1 | 63.8 |
@@ 3.4 Human Annotation
To support reliable automatic evaluation, we formalize the problem of complex commonsense reasoning as a multi-choice question answering task, with one true answer, three distractors, and a fifth option indicating “None of the answers are correct”. We crowdsourced the answers using Amazon Mechanical Turk (AMT). The workers are given the verbalized query as the context, the verbalized relations as the question, and the sampled (negative) answers. If no sampled answers are correct, then the worker is asked to select an additional “None of the answers are correct” option. If the verbalization itself does not make sense, the worker can also select another option “The context doesn’t make sense or is meaningless” and we discard the example. Each question is annotated by three workers. The overall per-option inter-annotator agreement is 78%, and the Fleiss kappa is 0.445, indicating moderate agreement. The workers are paid on average 16 USD per hour. Our final dataset consists of 782k training examples and 1317 manually-validated evaluation examples. More details can be found in Appendix A.
@@ 4 Experiments
We conduct experiments on the evaluation set of Com, formulated as a Multi-Choice Question Answering (MCQA) task. Specifically, we examine the performance of state-of-the-art off-the-shelf language models on Com, and also study the effect of training a question answering model on the distantly supervised training set of Com.
@@ 4.1 Setup
| Model | Training Data | Multi-Event | Paragraph-Level | Single-Event | Com | ||||||||
@@ 4.2 Results and Analysis
For models fine-tuned on complex queries using HyKAS and CAR, we observe that the synthetic training pairs, despite lacking manual annotation, serve as valuable distant supervision signals. They enhance the complex reasoning capability of HyKAS and CAR, surpassing the performance of the 8-shot GPT-4 model with CoT by 6%. CAR + Com also outperforms the 11B version of UnifiedQA-v2 and Flan-T5, which are both fine-tuned on numerous (commonsense) question answering datasets, by 9% and 3%, respectively.
@@ 5 Downstream Evaluation
In addition to benchmarking Complex Commonsense Reasoning, we also study the effect of leveraging Com as training data to generalize to other downstream commonsense reasoning tasks. As tasks, we use zero-shot CommonSense Question Answering (CSQA), and Generative Commonsense Inference, including one-hop, multi-event, and paragraph-level settings.
@@ 5.1 Commonsense Question Answering
The task of zero-shot commonsense QA involves selecting the most plausible option for commonsense questions without training on examples from the benchmark dataset. We directly leverage the model we trained in Section 4, the DeBERTa-v3-large-based model fine-tuned on synthetic question pairs from both ATOMIC and Com, and check the performance on five popular commonsense question answering datasets: Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). As baselines, we consider the same methods, HyKAS (Ma et al., 2021) and CAR (Wang et al., 2023a), but use other CSKGs as training sets. In Table 2, ATM-10X refers to ATOMIC-10x from West et al. (2022), and ATM refers to the training data from CAR Wang et al. (2023a) which is augmented from ATOMIC with conceptualization.
@@
We report model performance in Table 2. We observe the inclusion of Com and one-hop triples from ATOMIC as training data for CAR and HyKAS yields significant improvements in question answering ability. Notably, the combination of CAR and Com achieves the highest performance among all models, surpassing even ChatGPT and GPT-4, despite having a parameter size at least two orders of magnitude smaller.
Notably, when using CAR as the base model, training on Com leads to the highest performance gain of around 1.8% for a-NLI. When evaluating on a-NLI, which includes instances of abductive reasoning, the model may be helped by learning from 2i queries where one relation represents cause and the other represents effect (abduction examples in Figure 1 and Figure 3). Meanwhile, the performance on WinoGrande was adversely affected, likely because Winogrande primarily focuses on identifying distinguishing features of entity pairs. The benefits from learning event-event interactions from Com may not transfer well to this setting.
@@ 5.2 Generative Commonsense Inference
We study generative commonsense inference as an additional evaluation task. We include multi-event commonsense generation (COMET-M; Ravi et al., 2023) and paragraph-level commonsense generation (ParaCOMET; Gabriel et al., 2021) as two out-of-domain evaluation tasks. We also include the vanilla COMET Bosselut et al. (2019) as an additional in-domain evaluation, which focuses on 1p queries that require generating the tail given head and relation as the input. We also conduct experiments on the generative sub-task of Com, where verbalized context and question inputs are used to inferences. The annotated ground answer options are used as references.
For the (distantly) supervised learning baselines, we fine-tune GPT-2-large on the annotated multi-event inference dataset (MEI) from Ravi et al. (2023) and distantly labeled PCD dataset from Gabriel et al. (2021) as a reference. In our zero-shot learning setting, we study the effect of fine-tuning COMET (GPT-2-large) on ATOMIC and different query types of Com. We also study fine-tuning an LLM, Llama2-7b, by converting triples and queries to an instruction-tuning format, following the prompt template in Section 3.3 and Section B.2. We leverage the framework of Chen et al. (2023)333https://github.com/epfLLM to fine-tune Llama2-7b. We fine-tune on a mixture of different query types as detailed in the Training Data column. We present the performance results of models fine-tuned on either the annotated or distantly supervised training set for both tasks as reference benchmarks. Specifically, we use MEI for COMET-M and PCD for ParaCOMET. To ensure diversity and prevent overfitting to common tails, complex queries are selected using an n-gram based diversity filter Yang et al. (2020).
@@
6 Analysis & Discussion
6.1 Ablation Study
We analyze the impact of various data filters, query types, and verbalization methods on generative inference within Com. Detailed results can be found in Table 4.
Filtering
We include two types of filters, a Vera-based plausibility filter and a diversity filter. Evaluating the performance of generative commonsense inferences on Com, we examine the impact of removing both filters while employing GPT2-Large as the backbone model. Removing the plausibility filter results in a significant performance decline, highlighting its critical role. On the other hand, the diversity filter exhibits a minor positive influence on enhancing performance.
Type of Queries
| Model | Com | ||
@@ 6.1 Ablation Study
| Com-COMET | 13.6 | 26.1 | 39.8 |
| Com-COMET (V) | 14.3 | 27.1 | 43.4 |
| Com-Llama | 35.7 | 107.2 | 61.3 |
| Com-Llama (V) | 36.2 | 105.4 | 61.4 |
@@
| Com-COMET | 13.8 | 22.1 | 53.7 |
| Com-COMET (V) | 14.0 | 23.2 | 54.0 |
| Com-Llama | 14.6 | 22.1 | 55.3 |
| Com-Llama (V) | 14.8 | 23.6 | 55.5 |
@@
We investigate the effect of using a rule-based verbalizer or ChatGPT-enabled verbalizer to generate Com contexts. Using ChatGPT-verbalized queries leads to better downstream performance on both PCD and Com. In Com, the presence of ChatGPT-verbalization intuitively improves performance since the training context aligns with the evaluation set’s format. On the other hand, the context in the PCD dataset is long and comprised of five sentences. Verbalization not only adds more contexts to the training but also aligns better with the PCD format.
@@ 6.2 Error Analysis
We present a human-annotated quality evaluation of the Llama-7b-based model on the generation sub-task of Com. To ensure diverse coverage of query types, we randomly sampled 60 queries, with 10 from each of the 6 types. Manual inspection revealed a common error where the generated output was partially correct, either providing the answer to one of the triples in an intersection query or only the one-hop answer instead of the two-hop answer in 2-projection (2p) queries. Table 5 includes the number of such ‘1-hop’ partially correct answers. Our results demonstrate that the zero-shot Llama model already produces 26 out of 60 plausible inferences. Fine-tuning the model on one-hop ATOMIC further increases the number of plausible generations while more frequently generating inferences that are one-hop correct. Moreover, fine-tuning on the synthetic training set of Com significantly improves the model’s ability to generate complex commonsense inferences and reduces the occurrence of partially correct answers. We provide case studies in Appendix D.

### v2

@@ Abstract
Event commonsense reasoning requires the ability to reason about the relationship between events, as well as infer implicit context underlying that relationship. However, data scarcity makes it challenging for language models to learn to generate commonsense inferences for contexts and questions involving interactions between complex events. To address this demand, we present Com2 (COMplex COMmonsense), a new dataset created by sampling multi-hop logical queries (e.g., the joint effect or cause of both event A and B, or the effect of the effect of event C) from an existing commonsense knowledge graph (CSKG), and verbalizing them using handcrafted rules and large language models into multiple-choice and text generation questions.
Our experiments show that language models trained on Com2 exhibit significant improvements in complex reasoning ability, resulting in enhanced zero-shot performance in both in-domain and out-of-domain tasks for question answering and generative commonsense reasoning, without expensive human annotations.111Code and data are available at https://github.com/tqfang/complex-commonsense-reasoning
@@ 1 Introduction
In this paper, we construct Com2 (COMplex COMmonsense), a novel commonsense reasoning dataset using multi-hop queries in commonsense knowledge graphs to construct question answer pairs requiring complex narrative reasoning. To build this dataset, we use conjunctive logical queries Hamilton et al. (2018), a subset of First-Order Logical queries that use existential quantifiers and conjunction. The multi-hop projection operation involves inferring hidden contexts, while the intersection operation enables reasoning among multiple events, encompassing common cause or effect, and abduction. For example, in Figure 1, an intersection of two triples can be verbalized to a short narrative, and the process of inferring the common tail can be seen as an abduction of the hidden cause between the two heads.
To address the challenges above, we propose to first densify the CSKG to merge nodes with high semantic similarity, increasing the connectivity of the graph. Then, we use an off-the-shelf plausibility scorer to filter out low quality triples, avoiding error propagation as we construct more complicated queries. Finally, we verbalize the queries to a natural language context with handcrafted rules and Large Language Models to derive coherent and informative narrative contexts for our questions. Our final Com2 dataset comprises 790K question-answer pairs (both with multiple-choice and generative answer settings), including 1.3K examples that we manually verify for evaluation.
Our results demonstrate the challenges faced by even powerful LLMs and supervised question answering models on the Com2 dataset, underscoring the difficulty of performing complex multi-hop reasoning. Moreover, fine-tuning question answering models and generative commonsense inference models on Com2 leads to substantial improvements across eight commonsense reasoning datasets, showing the efficacy of our framework for boosting commonsense reasoning ability.
To conclude, our contributions are three-fold. First, we present a pipeline for sampling and verbalizing complex logical queries from CSKGs, to form a complex commonsense reasoning benchmark, Com2, with minimal human effort. Second, we benchmark the complex reasoning ability of various state-of-the-art language models and question answering models on Com2. Finally, we validate the benefit of fine-tuning on Com2 on eight zero-shot commonsense reasoning datasets.
@@ 2 Background and Related Work
Recent advances in commonsense reasoning have been driven by the construction of human-annotated (Speer et al., 2017; Sap et al., 2019a; Hwang et al., 2021; Jiang et al., 2021; Mostafazadeh et al., 2020; Krishna et al., 2017; Shen et al., 2024) and human-validated (West et al., 2022; Gao et al., 2023) CommonSense Knowledge Graphs (CSKG). A common approach to create challenges for commonsense reasoning involves constructing tasks in the form of question-answering Talmor et al. (2019); Sap et al. (2019b), knowledge base completion Malaviya et al. (2020); Yang et al. (2023) and population Fang et al. (2021b, a), grounding Gao et al. (2022), and daily dialogue Kim et al. (2023), based on CSKGs. However, most of those previous benchmarks are based on one-hop triples.
@@ 3 Methodology
In this section, we introduce the construction details of Com2, including the pre-processing, sampling, and verbalization of complex queries, as well as the details of human annotations. An ovewview of the pipeline is presented in Figure 2.
@@ 3.1 Pre-processing
CSKGs are usually highly sparse compared to factual KGs due to the diversity and scale of commonsense Malaviya et al. (2020), resulting in many isolated nodes that can hardly be sampled as part of a complex query. To alleviate this issue, we develop a set of rules and use sentence embedding similarity to merge nodes in the CSKG, leading to 22.4% of nodes being merged and an average degree increase of 25.3%. In the final query sampling process, the number of 2p paths increased from 7,382 to 405,492, and the number of 2i queries rose from 1.43M to 2.06M.
@@ 3.3 Verbalization
| Fine-tuned on Com2 | |||||||
| DeBERTa-v3-Large (+Com2) | 60.09 | 58.36 | 69.41 | 61.86 | 59.29 | 81.48 | 62.79 |
| CAR-DeBERTa-v3-Large (+Com2) | 61.22 | 56.13 | 69.86 | 68.56 | 56.64 | 85.19 | 63.78 |
@@
| a-NLI | CSQA | PIQA | SIQA | WG | Avg. | Com2 | ||
@@
| CAR Wang et al. (2023a) | ATMC | 79.6 | 69.3 | 78.6 | 64.0 | 78.2 | 73.9 | 39.8 |
| HyKAS + Com2(Ours) | ATM, Com2 | 78.4 | 69.9 | 78.7 | 64.1 | 78.3 | 73.9 | 62.8 |
| CAR + Com2(Ours) | ATM Com2 | 81.2 | 70.9 | 80.3 | 65.6 | 77.4 | 75.1 | 63.8 |
@@ 3.4 Human Annotation
To support reliable automatic evaluation, we formalize the problem of complex commonsense reasoning as a multi-choice question answering task, with one true answer, three distractors, and a fifth option indicating “None of the answers are correct”. We crowdsourced the answers using Amazon Mechanical Turk (AMT). The workers are given the verbalized query as the context, the verbalized relations as the question, and the sampled (negative) answers. If no sampled answers are correct, then the worker is asked to select an additional “None of the answers are correct” option. If the verbalization itself does not make sense, the worker can also select another option “The context doesn’t make sense or is meaningless” and we discard the example. Each question is annotated by three workers. The workers are paid on average 16 USD per hour. Our final dataset consists of 782k training examples and 1317 manually-validated evaluation examples.
Quality
The overall per-option inter-annotator agreement is 78%, and the Fleiss kappa is 0.445, indicating moderate agreement. Among 1.3K verified examples, 4.7% were labeled as incorrect contextualization. The likelihood that a sampled answer is the correct response to the contextualized question is 52.1%. For randomly sampled negative examples and one-hop neighbors, the plausibility rate is 23.5%, notably lower than the sampled answers. The authors of this paper manually checked the examples where the IAA between three annotators is lower than 0.6 and fixed the answers to ensure quality. A similar distribution is expected for the training set. Another thing to note that even though the training set is silver-standard, language models fine-tuned on it can autonomously identify patterns and acquire valuable insights from a large number of complex queries, resulting in improved reasoning performance, which will be shown in the next section.
More details can be found in Appendix A.
@@ 4 Experiments
We conduct experiments on the evaluation set of Com2, formulated as a Multi-Choice Question Answering (MCQA) task. Specifically, we examine the performance of state-of-the-art off-the-shelf language models on Com2, and also study the effect of training a question answering model on the distantly supervised training set of Com2.
@@ 4.1 Setup
| Model | Training Data | Multi-Event | Paragraph-Level | Single-Event | Com2 | ||||||||
@@ 4.2 Results and Analysis
For models fine-tuned on complex queries using HyKAS and CAR, we observe that the synthetic training pairs, despite lacking manual annotation, serve as valuable distant supervision signals. They enhance the complex reasoning capability of HyKAS and CAR, surpassing the performance of the 8-shot GPT-4 model with CoT by 6%. CAR + Com2 also outperforms the 11B version of UnifiedQA-v2 and Flan-T5, which are both fine-tuned on numerous (commonsense) question answering datasets, by 9% and 3%, respectively.
@@ 5 Downstream Evaluation
In addition to benchmarking Complex Commonsense Reasoning, we also study the effect of leveraging Com2 as training data to generalize to other downstream commonsense reasoning tasks. As tasks, we use zero-shot CommonSense Question Answering (CSQA), and Generative Commonsense Inference, including one-hop, multi-event, and paragraph-level settings.
@@ 5.1 Commonsense Question Answering
The task of zero-shot commonsense QA involves selecting the most plausible option for commonsense questions without training on examples from the benchmark dataset. We directly leverage the model we trained in Section 4, the DeBERTa-v3-large-based model fine-tuned on synthetic question pairs from both ATOMIC and Com2, and check the performance on five popular commonsense question answering datasets: Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). As baselines, we consider the same methods, HyKAS (Ma et al., 2021) and CAR (Wang et al., 2023a), but use other CSKGs as training sets. In Table 2, ATM-10X refers to ATOMIC-10x from West et al. (2022), and ATMC refers to the training data from CAR Wang et al. (2023a) which is augmented from ATOMIC with conceptualization.
@@
We report model performance in Table 2. We observe the inclusion of Com2 and one-hop triples from ATOMIC as training data for CAR and HyKAS yields significant improvements in question answering ability. Notably, the combination of CAR and Com2 achieves the highest performance among all models, surpassing even ChatGPT and GPT-4, despite having a parameter size at least two orders of magnitude smaller.
Notably, when using CAR as the base model, training on Com2 leads to the highest performance gain of around 1.8% for a-NLI. When evaluating on a-NLI, which includes instances of abductive reasoning, the model may be helped by learning from 2i queries where one relation represents cause and the other represents effect (abduction examples in Figure 1 and Figure 4). Meanwhile, the performance on WinoGrande was adversely affected, likely because Winogrande primarily focuses on identifying distinguishing features of entity pairs. The benefits from learning event-event interactions from Com2 may not transfer well to this setting.
@@ 5.2 Generative Commonsense Inference
We study generative commonsense inference as an additional evaluation task. We include multi-event commonsense generation (COMET-M; Ravi et al., 2023) and paragraph-level commonsense generation (ParaCOMET; Gabriel et al., 2021) as two out-of-domain evaluation tasks. We also include the vanilla COMET Bosselut et al. (2019) as an additional in-domain evaluation, which focuses on 1p queries that require generating the tail given head and relation as the input. We also conduct experiments on the generative sub-task of Com2, where verbalized context and question inputs are used to inferences. The annotated ground answer options are used as references.
For the (distantly) supervised learning baselines, we fine-tune GPT-2-large on the annotated multi-event inference dataset (MEI) from Ravi et al. (2023) and distantly labeled PCD dataset from Gabriel et al. (2021) as a reference. In our zero-shot learning setting, we study the effect of fine-tuning COMET (GPT-2-large) on ATOMIC and different query types of Com2. We also study fine-tuning an LLM, Llama2-7b, by converting triples and queries to an instruction-tuning format, following the prompt template in Section 3.3 and Section B.2. We leverage the framework of Chen et al. (2023)333https://github.com/epfLLM to fine-tune Llama2-7b. We fine-tune on a mixture of different query types as detailed in the Training Data column. We present the performance results of models fine-tuned on either the annotated or distantly supervised training set for both tasks as reference benchmarks. Specifically, we use MEI for COMET-M and PCD for ParaCOMET. To ensure diversity and prevent overfitting to common tails, complex queries are selected using an n-gram based diversity filter Yang et al. (2020).
@@
| Model | Com2 | ||
@@
| Com2-COMET | 13.6 | 26.1 | 39.8 |
| Com2-COMET (V) | 14.3 | 27.1 | 43.4 |
| Com2-Llama | 35.7 | 107.2 | 61.3 |
| Com2-Llama (V) | 36.2 | 105.4 | 61.4 |
@@
| Com2-COMET | 13.8 | 22.1 | 53.7 |
| Com2-COMET (V) | 14.0 | 23.2 | 54.0 |
| Com2-Llama | 14.6 | 22.1 | 55.3 |
| Com2-Llama (V) | 14.8 | 23.6 | 55.5 |
6 Analysis & Discussion
6.1 Ablation Study
We analyze the impact of various data filters, query types, and verbalization methods on generative inference within Com2. Detailed results can be found in Table 4.
Filtering
We include two types of filters, a Vera-based plausibility filter and a diversity filter. Evaluating the performance of generative commonsense inferences on Com2, we examine the impact of removing both filters while employing GPT2-Large as the backbone model. Removing the plausibility filter results in a significant performance decline, highlighting its critical role. On the other hand, the diversity filter exhibits a minor positive influence on enhancing performance.
Type of Queries
@@ 6.1 Ablation Study
We investigate the effect of using a rule-based verbalizer or ChatGPT-enabled verbalizer to generate Com2 contexts. Using ChatGPT-verbalized queries leads to better downstream performance on both PCD and Com2. In Com2, the presence of ChatGPT-verbalization intuitively improves performance since the training context aligns with the evaluation set’s format. On the other hand, the context in the PCD dataset is long and comprised of five sentences. Verbalization not only adds more contexts to the training but also aligns better with the PCD format.
@@ 6.2 Error Analysis
We present a human-annotated quality evaluation of the Llama-7b-based model on the generation sub-task of Com2. To ensure diverse coverage of query types, we randomly sampled 60 queries, with 10 from each of the 6 types. Manual inspection revealed a common error where the generated output was partially correct, either providing the answer to one of the triples in an intersection query or only the one-hop answer instead of the two-hop answer in 2-projection (2p) queries. Table 5 includes the number of such ‘1-hop’ partially correct answers. Our results demonstrate that the zero-shot Llama model already produces 26 out of 60 plausible inferences. Fine-tuning the model on one-hop ATOMIC further increases the number of plausible generations while more frequently generating inferences that are one-hop correct. Moreover, fine-tuning on the synthetic training set of Com2 significantly improves the model’s ability to generate complex commonsense inferences and reduces the occurrence of partially correct answers. We provide case studies in Appendix D.